OpenStudy (anonymous):

The area of a triangle is 30 square feet. If the height of the triangle is 5 feet, what is the base of the triangle?

OpenStudy (across):

\[A=\frac12bh\]

OpenStudy (anonymous):

area=0.5*base*height 30=0.5*5*b b=60/5 b=12

OpenStudy (anonymous):

base=12metres yar

OpenStudy (anonymous):

thank u @sai1234

OpenStudy (anonymous):

and across

OpenStudy (anonymous):

@across

OpenStudy (anonymous):

base is 3

OpenStudy (across):

How did you obtain that, @ArbY777?

OpenStudy (anonymous):

its 12 right

OpenStudy (anonymous):

\[A=\left(\begin{matrix}1 \\ 2\end{matrix}\right)b.h\]

OpenStudy (anonymous):

arb777 can u help me on my other problem

OpenStudy (anonymous):

triangle is a half of a square so the formula for triangle is the half of formula for sqaure so divide it by 2

OpenStudy (anonymous):

not sure if i can answer it all

OpenStudy (across):

@ArbY777, how did you conclude that the base is \(3\)? Also,\[\binom12\neq\frac{1}{2}.\]

OpenStudy (anonymous):

sory 12 is the answer ^^